OpenStudy (anonymous):

Suppose u(t)=w(t^2+1) and w′(2)=11 . Find u′(1) .

OpenStudy (goformit100):

Its a composite function... Simply Diffrentiare it...

OpenStudy (anonymous):

2t

OpenStudy (anonymous):

lets try this \[u(t)=w(t^2+1)\] \[u'(t)=w'(t^2+1)\times 2t\] by the chain rule so \[u'(1)=w'(1^2+1)\times 2\times 1=w'(2)\times 2=11\times 2=22\]

OpenStudy (anonymous):

thank you!